OpenStudy (dumbsearch2):

Which of the following is not a geometric sequence? A. 1, 2, 4, 8 B. 3, 1, -1,-3 C. ½, ¼, 1/8, 1/16 D. 216, 72, 24, 8

OpenStudy (dumbsearch2):

I thought it was D.

hartnn (hartnn):

it is

hartnn (hartnn):

and B too

hartnn (hartnn):

oh, infact its not D!

hartnn (hartnn):

there is a common ration in D, try diving next term to current term find common ratio

hartnn (hartnn):

216/72 =... 72/24 = ... 24/8 =...

OpenStudy (dumbsearch2):

Oh yes, I see it now! So it would be A.

OpenStudy (dumbsearch2):

No.

OpenStudy (dumbsearch2):

That's wrong.

hartnn (hartnn):

try B

OpenStudy (dumbsearch2):

It would be B. there is no real pattern :)

hartnn (hartnn):

correct :)